### v3

# From Large to Super-Tiny: End-to-End Optimization for Cost-Efficient LLMs

Abstract
Large Language Models (LLMs) have significantly advanced artificial intelligence by optimizing traditional Natural Language Processing (NLP) workflows, facilitating their integration into various systems. Many such NLP systems, including ours, directly incorporate LLMs. However, this approach either results in expensive costs or yields suboptimal performance after fine-tuning. In this paper, we introduce a three-stage cost-efficient end-to-end LLM deployment pipeline, comprising prototyping, knowledge transfer, and model compression, to effectively tackle the cost-performance dilemma in LLM-based frameworks. Its high cost-efficiency is manifested not only in simplifying system complexity and producing super-tiny online models with enhanced performance and reduced costs in the results, but also in addressing development cycle constraints, the lack of extensive high-quality data, and limited computational resources during the project development process. In the first stage, we construct an optimal performance prototype system by transforming complex tasks into a function call-based LLM-driven pipeline, which serves as a teacher model to generate high-quality data. In the second stage, we combine techniques like rejection sampling fine-tuning, reinforcement learning, and knowledge distillation to transfer knowledge to 0.5B student models, delivering effective performance at minimal cost. In the final stage, we further compress models to 0.4B via quantization and pruning, achieving ultra-low latency and cost. Extensive experimental results and the framework’s modular design suggest cross-domain capabilities and potential applicability in other NLP areas.
[…]
1 Introduction
In recent years, Large Language Models (LLMs) have made significant strides [1, 2, 3] in enhancing artificial intelligence by addressing shortcomings of traditional pipelines — including manual annotation, suboptimal performance, and limited generalization and adaptability. Consequently, LLM-based pipelines have stimulated substantial demand for integration into various Natural Language Processing (NLP) systems. We find that numerous such NLP systems, including our own, have directly utilized APIs of closed-source LLMs or fine-tuned open-source LLMs, which can be referred to as the "one-stage" pipeline. Although this pipeline offers implementation simplicity, it suffers from substantial costs [4] and high inference latency [5] or suboptimal performance, which limit the broader applicability of LLMs and impact their commercial viability. Within our NLP system, cost-performance is an especially critical metric, a sentiment likely shared by others in the field [6]. Moreover, our analysis indicates that this pipeline struggles to balance the cost and the performance due to insufficient high-quality data and monotonous optimization techniques.
In this paper, we introduce a three-stage cost-efficient end-to-end LLM deployment pipeline, as shown in Figure 1, consisting of prototyping, knowledge transfer, and model compression, to address the pervasive "cost-performance" dilemma in the practical deployment of LLM-based frameworks. The high cost-efficiency of this approach is demonstrated by both the outcomes and the process. In terms of outcomes, the approach simplifies system complexity and produces super-tiny models with optimal performance and reduced inference costs and latency. Regarding the process, the approach addresses challenges such as constrained development cycles, the scarcity of substantial high-quality data, and limited computational resources.
In our proposed pipeline, the first stage involves constructing the prototype system, where traditional complex tasks are converted into function call-based LLM-driven workflows [7] through prompt engineering, resulting in an optimal performance prototype compared to traditional pipeline. This prototype serves as a teacher model in the second stage, efficiently generating vast quantities of high-quality data. In the second stage, leveraging the optimal prototype from the first stage, we employ techniques including Rejection sampling Fine-Tuning (RFT), Reinforcement Learning (RL) [8, 9, 10], and multi-strategy knowledge distillation (KD) on the Qwen 2.5 series model [11]. Our hybrid strategy effectively transfers generalization capabilities and the abilities of generating high-quality CoT to a 0.5B student model. The derived student model is considerably smaller than the prototype model while achieving nearly comparable performance at minimal cost. In the third stage, we utilize quantization and pruning techniques to further compress the student model developed in the second phase, resulting in a remarkably compact 0.4B model. Ultimately, this super-tiny model is deployed in our online system. This model achieves hundreds-fold compression and extremely low latency and cost in exchange for a trivial and acceptable loss in performance. The modular design and superior cross-domain performance of the framework suggest its potential applicability in other NLP domains. Our major contributions are as follows:
  • The first validated three-stage cost-efficient end-to-end LLM pipeline, simplifying system complexity, achieves SOTA results and resolves the "cost-performance" dilemma. It realizes 180 times compression with nearly consistent performance compared to original LLMs and up to an absolute 14% performance increase over traditional BERT-based systems.
  • We propose the first hybrid knowledge transfer method by combining RL and KD. It effectively transfers the capabilities of LLMs to super-tiny models and achieves performance comparable and even superior to certain larger models on domain-specific tasks.
  • The proposed framework offers advantages such as low training, inference, and deployment costs, minimal project development expenses with fast cycles, and a strong potential for cross-domain transferability. It provides inspiration for developing standardized cross-domain solutions in both academia and industry, facilitating the widespread LLMs adoption.
Figure 1: Workflows of prototyping, knowledge transfer and compression.
2 Related Works
LLM for Function Calling & Evaluation
Leveraging in-context learning and reasoning capabilities, LLMs can select functions and parameters from provided lists based on user intent [12]. Fine-tuned 7B models like Gorilla [13] achieve GPT-4-level Function Call accuracy on datasets with thousands of APIs. Building on this, TinyAgent [14] enables function calling capabilities for edge devices with sub-1.1B parameter models. Query-to-API reformulation via rejection sampling can also optimize API retrieval, reducing retrieval latency while improving accuracy [15]. Nowadays, LLMs can be used to evaluate model-generated content [16, 17]. CriticGPT [18] indicates that LLMs can build a cost-effective and high-precision judge, surpassing humans in detecting subtle errors.
Knowledge Distillation
Knowledge distillation (KD) [19] transfers knowledge from a teacher model to a smaller student model by simulating probability distribution. Common approaches include Logits-based[20, 21], Feature-based [22] and sequence-level (Seq-KD) [23] distillation. Initially, Logits-based KD uses the forward KL divergence (FKL) [24, 25] as the loss function. Some KD methods prefer Reverse KL divergence (RKL) [26, 27, 28, 21]. Recently, Adaptive Kullback-Leiber (AKL) divergence [29] achieves a balance between mimicking distribution and avoiding overfitting.
LLM Pruning & Quantization
Efficient LLM deployment primarily relies on pruning[30, 31] and quantization[32, 33, 34]. Structured pruning[35] employs layer removal[36, 37] or dimensionality reduction[31, 37, 38] using gradient-based[31, 38] or activation-aware[37] importance criteria, implemented via learnable masks[38, 39] or pruning-retraining pipelines[36, 31, 37]. Weight-only LLM quantization methods[32, 33, 40] improve memory efficiency while joint weight-activation schemes[34, 41] and their combination [42] accelerate computation.
Reinforcement Learning
Recent works [8, 9, 10] focus on leveraging Reinforcement Learning (RL) to enhance model reasoning capabilities. Notable reasoning models such as DeepSeek-R1 [43], QWQ-32B [44], and OpenAI o1 [45] have emerged, employing RL algorithms like Proximal Policy Optimization (PPO) [46], Group Relative Policy Optimization (GRPO) [47], and others [48, 49, 50]. For general-purpose LLMs, preference optimization algorithms [51, 52, 53, 54] are widely adopted. Additionally, self-improving pipelines [55, 56, 57] have been proposed to iteratively refine model performance. We primarily utilize RL to enhance the model’s generalization. Compared to directly distilling a reasoning model, RL results in shorter and more precise Chains-of-Thought (CoT).
3 Methodology
Recent advancements in LLMs have facilitated significant progress in NLP tasks such as classification and summarization. We harness these capabilities to reconstruct the traditional dialogue system through a cost-efficient end-to-end LLM pipeline, which includes prototyping, knowledge transfer, and model compression, as illustrated in 1(a) and 1(b). Initially, we develop the prototype system by leveraging the exceptional performance of contemporary LLMs. Subsequently, we perform knowledge transfer from large models to smaller models. Finally, we achieve additional reduction in model size through further model compression techniques. Ultimately, our objective is to deliver intelligent user experiences at minimal costs, as demonstrated in 1(c).
3.1 Prototyping
LLM Prototype
Traditional dialogue systems usually employ the Intent Classification and Slot Filling (ICSF) framework, often utilizing BERT[58, 59]. Given the limitations of a single model in handling complex real-world issues, a Domain Classifier (DC) is often utilized for initial query categorization to facilitate domain-specific ICSF parsing. Although the hierarchical DC-ICSF architecture improves accuracy, the system complexity increases with feature updates. This leads to high inter-module coupling, increasing development and maintenance costs. Furthermore, ambiguous domain boundaries turn the DC into a bottleneck for overall end-to-end accuracy, creating a "feature expansion - operational difficulty" feedback loop that hinders dialogue experience optimization.
Modern LLMs exhibit adequate capability to tackle these NLP problems. Particularly in related Tool Use works, LLMs can learn to call functions without additional training by attaching function descriptions in prompts. Essentially, ICSF can be viewed as a function call task, where the model extracts critical information from user queries as function parameters for specific modules to execute. Following this approach, we design a prototype system based on Qwen2.5-72B-Instruct[11]. As illustrated in 1(a), the prototype system accepts user queries and contextual information (including chat history and temporal status information) as input, searches relevant APIs from the API pool, then outputs function names and parameters to call. The traditional system comprises one DC module and multiple ICSF models (based on BERT or regular expressions), each requiring separate maintenance and frequent DC updates. In contrast, the LLM-based system, equipped with function call capability, processes all tasks through a unified model, thereby reducing system complexity and enhancing operational efficiency.
To enhance temporal awareness and mitigate hallucinations, we enrich the input with contextual knowledge via RAG [60], retrieving explanations and extensions of entity-related terms from a curated knowledge base. Instead of adding all callable APIs from the API pool into the LLM prompt, we implement an API retrieval module inspired by [15]. The LLM first generates potential API descriptions, then retrieves top-K APIs via vector similarity (using GTE-Qwen2-1.5B [61] as the embedding model and FAISS [62] for indexing). Then, the LLM selects optimal functions/parameters from candidate APIs. For backward compatibility, we map functions/parameters to legacy domain-intent-slot structures.
Judge
Traditional ICSF evaluation relies on test sets demanding strict intent-slot alignment, contrasting with real-world variability where utterances permit multiple valid interpretations. For example, "Slightly increase the volume" allows alternative API calls, and "11:00 pm"/"23:00" are semantically equivalent. This inherent ambiguity and complex business logic significantly impede consistent human annotation for accurate ground truth.
To address issues above, we introduce LLM-as-a-Judge to evaluate prototype accuracy. We design a judge based on QwQ-32B[44]: Given context and a set of Function & parameters, the judge determines whether this combination semantically satisfies the user query. Following Samuylova’s best practices [^1] https://www.evidentlyai.com/llm-guide/llm-as-a-judge, we optimize the judge using a 1k-expert-annotated dataset. Subsequently, both the LLM and 9 domain experts evaluate 3k test samples under identical conditions. Results show that judge can provide logically coherent Chain-of-Thought(CoT) rationales and it achieves 93% consistency with expert judgments. Based on these findings, we conclude that LLM-as-a-Judge is the optimal evaluation approach for this task. We reconstruct the dialogue pipeline from BERT-based DC-ICSF architecture to LLM-based Function Call architecture. Evaluation through LLM-as-a-Judge demonstrates that the new architecture improves accuracy from 81.43% to 94.74% compared to the original one.
3.2 Knowledge Transfer
Rejection sampling Fine-Tuning
Building on our agent-oriented large model design, we propose Supervised Fine-Tuning (SFT) on smaller models using large model-generated data instead of relying solely on in-context learning (ICL) with large models. This approach offers two key benefits: (1) Smaller models can achieve performance comparable to large models after trained on sufficient high-quality data; (2) It significantly boosts inference and optimization efficiency. For a dataset $D$ containing prompt component $x$ and response component $y$, we define $D(x,y)=\{x_{i},y_{i}\}_{i=1}^{N}$, where $N$ denotes the dataset size.
SFT employs Cross Entropy (CE) loss [63] as the optimization objective. Building on this, we apply Rejection sampling Fine-Tuning (RFT), utilizing an LLM judge, as introduced in Section 3.1, to assess training data generated by the LLM prototype, filter out low-quality data, and construct the final training dataset. The actual training data is formulated as:
$D_{filter}=\{x_{i},y_{i}\lvert\ \text{judge\_data}(x_{i},y_{i})=1\}_{i=1}^{N}$ (1)
where
[…]
Knowledge Distillation
Knowledge Distillation (KD) [19] offers an alternative to RFT for transferring capabilities from large to small models, typically requiring fewer data samples and exhibiting lower overfitting risks. We adopt the logits-based Adaptive Kullback-Leibler (AKL) distillation [29]:
$L_{AKL}=\mathbb{E}_{(x,y)\sim D}[AKL(\pi_{\phi}(y\lvert x),\pi_{\theta}(y\lvert x))]$ (3)
$AKL(p,q)=\alpha_{head}KL(p,q)+(1-\alpha_{head})KL(q,p)$ (4)
$KL(p,q)=\sum_{t}{p_{t}}\log\frac{p_{t}}{q_{t}}$ (5)
where $\pi_{\theta}$ denotes the trainable student model, $\pi_{\phi}$ represents the teacher model, and $\alpha_{head}$ is the head ratio in AKL. Setting $\alpha_{head}$ = 1 yields Forward KL Divergence-based distillation [25], while $\alpha_{head}$ = 0 corresponds to Reverse KL Divergence-based distillation [21].
We make optimizations on original methods: (1) Teacher Calibration: Following [64], we calibrate the teacher model with task-specific data to achieve lower student model loss; (2) Logits Pre-storage: For reducing training time, we pre-store teacher logits to enable multi-student distillation from a single cached logits set; (3) Top-k Pruning: We trim low-probability logits through top-k pruning, mitigating storage pressure and accelerating training speed.
Hybrid Knowledge Transfer
Figure 2: The workflow of the hybrid knowledge transfer strategy.
Following the standard knowledge transfer strategy, we employ either RFT or KD combined with prompt compression, which removes unnecessary format specifications and Chain-of-Thought (CoT) rationales (see Appendix B), to create a fast but domain-specific model. However, our experiments reveal limitations of this conventional approach. Utilizing RFT or KD alone can improve inference efficiency but fails to enhance out-of-domain generalization. According to Logic-RL [9] with task-specific reward adjustment (see Appendix A), we employ Reinforcement Learning (RL) to improve generalization. Additional experiments indicate that employing RL alone generally leads to varying effectiveness across different model sizes: RL training yields exceptional results on small models (7B), surpassing both the LLM prototype and RFT-trained models. On the other hand, when applied to super-tiny models (0.5B), RL becomes ineffective as these models tend to generate shorter or meaningless CoT responses, leading to degraded output diversity.
These findings motivate us to propose a hybrid knowledge transfer strategy, as illustrated in Figure 2. This hybrid strategy first employs RL to train models that are as small as possible while ensuring robust generalization and the ability to generate high-quality CoT, exemplified by the 7B RL-trained model in our experiments. Subsequently, KD is used to transfer these generalization capabilities and CoT generation abilities to super-tiny models. With CoT generation abilities, super-tiny models can be trained through a second round of RL to further enhance generalization. Experimental results demonstrate that super-tiny models exhibit impressive improvement in both in-domain and out-of-domain performance after the second RL training. Detailed performance comparison of different transfer strategies shows in Appendix J.
3.3 Model Compression
Pruning
To further compress model, we design two phases of pruning following NVIDIA [64], Llama3.2 [65], and Apple [66]: (1) Offline One-Shot Struct Pruning: Following Minitron[31], we explore two pruning strategies. For depth pruning, we compute layer importance scores and remove the $N$ least important layers based on the target pruning ratio. For width pruning, we conduct a model architecture search using PPL-based criteria and prune channels according to group-wise importance rankings, with full implementation details in Appendix H; (2) Recovery training: The first step is General Instruction Fine-Tuning via Knowledge Distillation. Adopting Minitron’s approach, we leverage logits from a larger general instruction teacher model for recovery training. Due to unavailability of the teacher’s original training data, we first perform model correction on our selected IFT dataset to stabilize distillation, followed by top-k logits distillation(Section 3.2). The second step is Domain-Specific Task Alignment. The pruned compact LM undergoes further aligning through knowledge distillation, producing a final domain-optimized LLM.
Quantization
Quantization typically reduces GPU memory consumption and enhances inference performance. In this work, we apply post-training quantization to smaller models obtained through distillation/pruning. Specifically, we implement two quantization approaches: 4-bit weight-only GPTQ algorithm [33] and full 8-bit weight-activation FP8 quantization [41]. The 4-bit weight-only quantization [33] compresses linear weights from FP16 to per-group INT4 via integer quantization, which are dequantized to FP16 during computation. This approach achieves notable inference acceleration in memory-bound scenarios with large model parameters and low GPU memory bandwidth. The FP8 quantization [41] quantizes both weights and activations to 8-bit floating-point format. Compared to 8-bit integer quantization [41], FP8 shows a more noticeable improvement in accuracy, and by utilizing NVIDIA’s FP8 tensor cores, FP8 can achieve effective inference acceleration [67].
4 Experiments
4.1 Experimental Setup
Datasets
In our experiments, we utilize a large-scale dataset built from years of team experience and practice accumulated in the field. We reconstruct the intent classification and semantic slot framework within the real-world system into an API document specification, creating a function pool comprising 74 functional modules with 225 arguments. We sample 200k data points from a single business scenario as the training set and synthesize labels based on the LLM prototype outlined in Section 3.1. A combination of stratified and random sampling yields 11,922 samples from this scenario’s data as the in-domain test set. Additionally, we sample 7,867 data points using the same methodology from another business scenario, without some callable functions in the training set, as the out-of-domain test set.
Metrics
In accordance with the LLM-based judge outlined in Section 3.1, we establish our evaluation metrics. Consider a dataset of size $N$, denoted as $D(x,y)=\{x_{i},y_{i}\}_{i=1}^{N}$, where $x$ signifies the prompt, including elements such as user queries, function lists, and dialogue histories, while $y$ represents the response containing the function and its parameters. The function judge_data, as rigorously defined in Equation 2, assesses whether the function and its parameter list can fulfill the specified user query. We define the Achievable Rate (AR) as follows:
[…]
Training
During the training phases, base models are all from small instruct models (0.5B-7B) of the Qwen2.5 model family [11]. We conduct RFT on them and RL on 7B models. We perform pruning on 0.5B models and obtain Qwen2.5-0.4B-Width and Qwen2.5-0.4B-Depth, both under a 20% pruning ratio. Then, we transfer knowledge from 7B RFT-trained models to pruned and 0.5B models via KD. We also apply KD from 7B RL-trained models to 0.5B models, followed by another RL phase (denoted as KD*+RL). Finally, we use model quantization via W8A16/W4A16/W8A8 strategies on the KD-trained 0.5B model, detailed in Appendix D. The training process is cost-efficient, utilizing only 8 NVIDIA H20 GPUs. Computing resources are listed in Appendix E.
Evaluation settings
We do evaluation by comparing our models with the BERT-based system consisting of two models based on BERT-large (0.33B)[68]. To evaluate the effectiveness of our method, we invite expert annotators alongside the LLM-as-a-Judge-based metric AR (see Equation 6) for performance assessment. We also implement a retrieval module to select top 10 functions from the function pool, aiming to accelerate both training and inference phases, as detailed in Appendix C. To further verify the effectiveness of our proposed approach, we train models on the xlam-function-calling-60k dataset generated by APIGen[69] and conduct validation using the Berkeley Function-Calling Leaderboard (BFCL)[70]. Details can be found in Appendix K. In real-world scenarios, we deploy models on NVIDIA L20 GPU and perform effectiveness evaluation via TensorRT-LLM[67], with specific configuration details provided in Appendix G. During the comparison between the trained models and the BERT-based system, we adopt queries per second (QPS) as the primary metric for system throughput, response time (RT) for end-to-end latency, and concurrent clients (CC) to quantify high-concurrency performance.
4.2 Experimental Results
Table 1: Main results of the models trained on Qwen2.5 (See Appendix I for FLOPs analysis).
Model / System AR FLOPs (G)
BERT-based system 81.43% 38.65
Qwen2.5-72B (ICL) 94.74% 4907.50
Qwen2.5-7B (RFT) 95.29% 434.04
Qwen2.5-3B (RFT) 95.30% 218.94
Qwen2.5-1.5B (RFT) 95.16% 111.91
Qwen2.5-0.5B (RFT) 94.66% 43.31
Qwen2.5-0.5B (KD) 94.85% 43.31
Qwen2.5-0.5B-W8A16 (KD) 94.76% -
Qwen2.5-0.5B-W4A16 (KD) 94.19% -
Qwen2.5-0.5B-W8A8 (KD) 94.83% -
Qwen2.5-0.4B-Width (KD) 94.17% 40.83
Qwen2.5-0.4B-Depth (KD) 94.20% 33.10
Figure 3: Cost and performance results. FLOPs reduced by 10x with nearly lossless AR.
Table 2: In-domain and out-of-domain performances of different knowledge transfer strategies.
Model AR
AR
Qwen2.5-72B (ICL) 94.74% 94.63%
Qwen2.5-7B (RFT) 95.29% 90.86%
Qwen2.5-0.5B (RFT) 94.66% 90.62%
Qwen2.5-7B (RL) 95.36% 94.80%
Qwen2.5-0.5B (KD*+RL) 95.99% 92.34%
Figure 4: Comparison of annotation results between the LLM prototype and the BERT-based system determined by professional annotators.
LLM Prototype Preferred by Both Human and LLM-as-a-Judge
As illustrated in Figure 4, annotations provided by professional annotators reveal that the LLM prototype (72B ICL) achieves a superior performance over the BERT-based system, with a win+tie rate of 95.59%. Through the evaluation via LLM-as-a-Judge, we find that the LLM Prototype surpasses the BERT-based system by 13.31%, as shown in Table 1. This conclusion is nearly consistent with that presented by professional annotators, demonstrating the effectiveness of the LLM judge.
RFT and KD as Effective Knowledge Transfer Methods
As demonstrated in Table 1, models trained with RFT generally exhibit a higher AR than the LLM prototype. Specifically, both the 3B and 7B models display negligible differences in performance and both achieves a 0.5% improvement in AR compared to the LLM prototype. The 0.5B model’s AR is 0.5% lower than the 1.5B model, while KD emerges as a valuable complement to RFT, enhancing the performance of the 0.5B model by an additional 0.19%. In terms of inference performance, the 0.5B model demonstrates only a slight 0.4% difference in AR compared to the 7B model, while achieving a significant 13.57x improvement in QPS and a substantial 91.3% reduction in latency, as illustrated in Table 3.
RL’s Superior Generalization Ability
According to Table 2, while RFT performs well on in-domain data, it demonstrates worse results on out-of-domain data with a 3.61% AR decrease compared to the LLM prototype. Conversely, 7B RL-trained models achieve superior performance compared to other models on both in-domain and out-of-domain datasets, indicating better generalization of RL training methodology. Notably, even with models as small as 0.5B, employing a hybrid training methodology with KD+RL can yield superior in-domain results and an improvement of 1.7% in out-of-domain datasets compared to the RFT method. For more detailed exploration of the RL results, please refer to Appendix J.
Tiny LLMs as a Cost-Efficient Upgrade of BERT-based System
As shown in Table 1, Table 3 and Figure 5, quantization minimally impacts accuracy, with less than a 0.1% AR loss, while significantly enhancing inference performance, achieving up to 1.29x higher QPS and 0.76x lower RT in the optimal scenario. Pruning incurs a slightly higher 0.2% accuracy loss but consistently enhances performance by 20%. For end-to-end inference, with a CC of 64, the BERT-based system reaches 1110.88 QPS and 55.84 ms RT. Although the cost is roughly twice that of the BERT-based system, the AR improves by up to absolute 14%.
Table 3: Comparison of QPS metrics using different quantization and pruning strategies.
Model QPS
cc=1 cc=8 cc=16 cc=32 cc=64 cc=128
Qwen2.5-0.5B 39.11 166.71 248.62 348.87 459.12 517.13
Qwen2.5-0.5B-W8A16 44.11 169.86 251.21 346.17 445.84 521.41
Qwen2.5-0.5B-W4A16 49.83 175.29 256.05 338.16 453.46 519.22
Qwen2.5-0.5B-W8A8 40.40 158.15 256.32 382.13 558.53 668.35
Qwen2.5-0.4B-Width 43.65 167.65 260.58 372.59 502.82 581.23
Qwen2.5-0.4B-Depth 45.22 198.82 309.62 438.58 566.87 650.26
Qwen2.5-7B 4.43 14.77 17.71 19.04 19.79 20.06
Figure 5: Comparison of RT metrics using different quantization and pruning strategies.
5 Analysis
Performance-Efficiency Trade-offs in Model Scaling
Our experimental analysis highlights a compelling trade-off between model size and computational efficiency. Although larger models typically show superior performance, the expected disparity between large and small ones is not as pronounced as anticipated. As shown in Figure 3, the 0.4B model achieves performance almost on par with the 7B model, efficiently requiring 11.8% FLOPs and reaching more than 17.4x QPS improvement and 92.9% RT reduction, with the 7B model’s RT being 882.552 ms under 16 CC. Although this model has a slight gap in cost compared to traditional BERT models, the improved accuracy and scalability stemming from LLM base makes it a favorable choice for various applications.
Characteristics of Training Paradigms
The results presented in Table 2 underscore the distinct characteristics of different training methodologies. Fine-tuning primarily focuses on memorizing data, making it susceptible to overfitting to the training set. In contrast, RL presents greater robustness by not merely fitting individual data points. We posit that RL’s superior generalization capabilities stem from its ability to learn from negative samples, which are generated by the model being trained itself. RL is highly effective for dealing with well-structured problems, while it has unsatisfactory convergence facing ambiguous data or a lack of domain knowledge in real-world tasks. Additionally, applying RL directly to tiny models often proves to be ineffective, likely due to such models’ limited capacity to utilize Chain-of-Thought (CoT) rationales. In such cases, KD can compensate for this limitation by training the tiny models on a large amount of CoT data generated by a RL-trained teacher model. Compared to fine-tuning, KD can better extract CoT rationales from the probability distribution. Consequently, the hybrid methodology of KD and RL can achieve remarkably enhanced performance and represents a promising direction for future development.
Comparison over Model Compression Strategies
Our analysis identifies depth pruning as the optimal approach for domain-specific models. Depth pruning achieves 20% inference acceleration with lossless AR at same compression ratios (Table 1), stemming from the nature of depth pruning, removing entire transformer layers and their associated attention computations. Although width pruning exhibits slightly better effects in general instruction tasks after training (Table 8), depth pruning presents optimal efficiency facing long prompts, with width pruning focusing on matrix optimizations (Appendix I). Regarding quantization, the W4A16 strategy achieves impressive performance in single-thread scenarios, while FP8 outperforms with lossless accuracy under high concurrency conditions as the superior paradigm for throughput deployments.
6 Conclusion
Based on our comprehensive study, several key insights emerge regarding the development and deployment of effective LLMs in practical scenarios.
A Scalable Practical Alternative for Cost-Efficient LLMs
Our proposed pipeline presents a systematic framework designed to achieve high performance while maintaining efficient costs in both development and deployment. Our end-to-end optimized tiny models also outperform their larger counterparts, offering resource efficiency without compromising accuracy. This provides an alternative pathway for applying LLMs in practical tasks under low-cost scenarios. We believe this approach can be effectively adapted to other domains, thereby potentially advancing research and facilitating the widespread deployment of LLMs.
LLM Capability Preservation via a Hybrid Strategy
Our hybrid strategy demonstrates strong effectiveness and practicality in preserving LLMs’ capabilities during knowledge transfer. It enhances generalization and high-quality CoT generation abilities via RL, safely retains these capabilities in smaller models via KD, and further improves the in-domain and out-of-domain performance of super-tiny models via another RL. This methodology judiciously and scientifically avoids irreversible loss of capabilities that might occur when directly deriving super-tiny models via RL. It progressively retains LLMs’ capabilities, transferring them step by step to super-tiny models.
Balanced Focus Beyond AGI
While AGI remains a visionary long-term aim for LLMs, progress has reached a plateau, prompting increased attention towards real-world applicability. Our work demonstrates that integrating LLM-driven techniques into traditional tasks results in immediate, measurable improvements. It advocates for a balanced research agenda that equally emphasizes foundational advancements and domain-specific applications.
7 Limitation & Future Works
Our pipeline has presented optimal cost-performance results, and our super-tiny models have exhibited remarkable capabilities, featuring low latency, cost efficiency, cloud independence, and robust privacy. Further compression may offer limited marginal benefits. Thus, we are more focused on enhancing such models’ capabilities and expand their broader applications.
Deeper Study on LLM Capability Transfer
Based on our experiments, the hybrid strategy effectively enhances super-tiny models’ generalization while almost entirely retaining LLMs’ capabilities in domain-specific tasks. In future research, we will conduct a deeper investigation and theoretical analysis of this potentially powerful methodology, building upon prior work to gain a better understanding of this hybrid strategy in transferring LLMs’ capabilities to super-tiny models.
Cross-domain Optimal Compression Exploration
It is evident that compressing models to 0.5B while retaining their capabilities across all domains is not feasible. The limitation boundary remains unclear. So, further research and exploration are required to better understand and define these limits.
Appendix A Details for Reinforcement Learning
Our reward design consists of two components, Format Reward($S_{format}$) and Answer Reward ($S_{answer}$):
  • For Format Reward, we adopt the logic of Logic-RL[9], which validates the output format against the specified template <think>...</think><answer>...</answer> via regular expressions. We add additional designs to ensure the validity of the usage of the function and its parameters. If a function not present in the predefined input function pool is selected, it is deemed invalid.
$S_{format}=\begin{cases}1,&\text{if the format is correct, and the function call is valid}\\ 0,&\text{otherwise}\end{cases}$ (7)
  • For Answer Reward, in our specific scenario, we observe that the model tends to directly guess answer during training, resulting in no association between the Chain-of-Thought (CoT) and the final results. Requiring explicit repetition of the answer within the CoT framework effectively resolves this issue. As illustrated in Figure 6, we optimize its original design in order to enhance the consistency between the Chain-of-Thought (CoT) reasoning and the results. Our requirements are as follows: (1) The CoT must explicitly reference the target function at least once; (2) The final function mentioned in the CoT must correspond to the function in the answer. Failure to comply with these conditions will be regarded as a parsing error. Additionally, we incorporate fine-grained rewards, which provide better guidance when the answer is partially correct, enabling the model to achieve improved performance.
$S_{answer}=\begin{cases}-2&\text{if the answer is missing, or is unable to be parsed,}\\ &\text{or is not defined in the function list,}\\ &\text{or the thought doesn't end with the function in the answer}\\ -1.25&\text{if the names of at least one parameter and function are correct,}\\ -1.2&\text{if the name of function, as well as the name }\\ &\text{and value of at least one parameter are correct}\\ 2,&\text{if the answer is correct,}\\ -1.5,&\text{otherwise}\end{cases}$ (8)
Figure 6: An example of optimization of the original Answer Reward design by requiring explicit repetition of the answer within the CoT framework.
Appendix B Prompt Compression
To reduce inference costs, we remove unnecessary content in prompts, such as guidelines, format specifications, and few-shot examples used during large model invocation, while retaining essential elements: user queries, context, query-relevant knowledge, function lists, and brief function descriptions. For outputs, we remove Chain-of-Thought (CoT) components and directly use predefined JSON-formatted responses, as illustrated in Figure 7.
Figure 7: Original vs Compressed Prompt & Response.
Appendix C Retriever Setup
Initially, we select GTE-large [61], which is trained on the BERT-large [68], as our embedding model. Using this model, we generate embeddings of API descriptions produced by the LLMs. Then, we compare these embeddings with those of the API description from the document using cosine similarity. However, this model’s recall accuracy is not high. Subsequently, we switch to GTE-Qwen2-1.5B-instruct [61], a model trained on Qwen2-1.5B [71], which offers the advantage of fine-tuning for matching tasks through prompt modification. Next, utilizing this model with a customized prompt in Section C.1, we perform cosine similarity matching between the model-generated API descriptions and the JSON text of the entire API documentation, achieving a higher accuracy. During training, we optimize the API description model using rejection sampling, with implementation details provided in Section C.2. Finally, to achieve optimal inference acceleration, we train a classifier based on BERT-large [68] for top-1 recall using rejection sampling. The classifier takes Dialog, User Query, and necessary context as input and the function category as output. The functions corresponding to the top-k output logits are then considered as candidate functions for the subsequent model’s function call. This approach significantly accelerates inference due to the efficiency of BERT classifiers compared to autoregressive GPT models, while also enhancing function recall capabilities. However, it introduces scalability challenges since the recall model needs to be retrained with each update to the function pool. As illustrated in Table 4, this series of optimizations lead to a substantially improvement in API recall rates.
Table 4: Comparison of retrieval performance of different models: GTE-large model, GTE-Qwen2-1.5B-instruct model, optimized API description model + GTE-Qwen2-1.5B-instruct model. Extensibility implies that the API pool can be dynamically modified.
Top-N GTE-large GTE-Qwen2 Opt-API+GTE-Qwen2 RFT-BERT
5 60.19% 88.27% 93.39% 99.61%
7 74.06% 94.67% 98.09% 99.92%
10 90.11% 98.22% 99.55% 99.92%
15 98.95% 99.80% 100% 100%
Extensibility ×
C.1 Prompt for retrieval
{mdframed}
Instruction: Given a API description, retrieve the most similar API from the function pool.
Query:
C.2 Retriever Optimization
Algorithm 1 Simplified API Description Validation Process
0:  Query $q$, LLM $\mathcal{M}$, API Pool $\mathcal{R}$, Retrieval Module $\mathcal{S}$, Function Call Module $\mathcal{F}$, Judge $\mathcal{J}$
0:  High-quality training dataset $\mathcal{D}$
1:  Generate multiple API descriptions $Des=\{\mathcal{M}(q)\}$ in parallel.
2:  for each description $d\in Des$ do
3:     Retrieve candidate APIs $C=\mathcal{S}(d,\mathcal{R})$.
4:     Select API $a\in C$ and generate parameters $p$.
5:     Execute $a$ with $p$, get result $r$.
6:     if $\mathcal{J}(r)=\text{Success}$ then
7:        Check if $a$ is in the top 5 candidates of $C$.
8:        if Yes then
9:           Add $(q,d,a,p,r)$ to $\mathcal{D}$.
10:        end if
11:     end if
12:  end for
Appendix D Training Details
Table 5: Training parameters of different knowledge transfer strategies.
Supervised Fine-Tuning (SFT)
Learning Rate $2\times 10^{-5}$
Batch Size $128$
Number of Epochs $1$
Weight Decay $1\times 10^{-4}$
Reinforcement Learning (RL)
Learning Rate $3\times 10^{-7}$
Batch Size $128$
Number of Episodes $1400$
Number of Rollouts $8$
KL Coefficient $0.001$
Knowledge Distillation (KD)
Learning Rate $2\times 10^{-5}$
Batch Size $128$
Number of Epochs $1$
Logits Top-K $100$
Table 6: Ablation of KD on Qwen2.5-0.5B models.
Model AR
Qwen2.5-0.5B (RFT) 94.66%
Qwen2.5-0.5B (FKL) 94.65%
Qwen2.5-0.5B (AKL[$\mu=0.5$]) 94.48%
Qwen2.5-0.5B (AKL[$\mu=0.7$]) 94.68%
Qwen2.5-0.5B (AKL[$\mu=0.9$]) 94.85%
The main training parameters are detailed in Table 6. During the training phase, we employ the AdamW optimizer [72] with the cosine learning rate schedule, setting the model’s maximum sequence length to $5120$. In the RL phase, we utilize the baseline variant of Reinforce++ [50] as the reinforcement learning algorithm, offering advantages in efficiency, conciseness, and training stability. Following [73], the K2 estimator is used instead of the original K1 in Reinforce++. In the AKL phase, the threshold parameter $\mu$ is reduced from 1 to 0.5. When training with data containing CoT, we set the number of training epochs to 2 and use the average of the FKL loss and the CE loss. For pruning, we use the 0.5B-Instruct model as the base, implementing width/depth pruning strategies with a 20%/30% prune ratio, resulting in two models Qwen2.5-0.4B(0.35B)-Width and Qwen2.5-0.4B(0.35B)-Depth. These extremely compressed models are produced through a two-phase training process involving accuracy recovery training and knowledge distillation (KD). Comprehensive training details for the pruned models are available in Appendix H. For quantization, we primarily experiment with three strategies on the 0.5B model trained with AKL: W8A16, W4A16, and W8A8. Both W8A16 and W4A16 employ per-group symmetric int8 weight-only quantization with a group_size=128, while activations are in fp16. To improve accuracy, W4A16 use 512 training data samples combined with the GPTQ algorithm [33] for calibration. In contrast, W8A8 uses FP8 for both weights and activations. For the training frameworks, we utilize Megatron-LM [74] for SFT and KD, while OpenRLHF [75] is employed for RL. Data sampling during training is conducted using vLLM [76]. In the evaluation phase, we employ SGLang [77] to generate data.
Appendix E Computing resources
We utilize 8 NVIDIA H20 GPUs for training and data generation processes. For evaluating inference performance, a single NVIDIA L20 GPU is employed. We apply an INTEL XEON PLATINUM 8575C processor with 192 cores for computation, operating at a frequency of 3028.956 MHz and a cache size of 327680 KB. Additionally, we establish a virtual development environment based on Kubernetes, leveraging 64 virtual cores and 512 GB of memory. During the training phase, the RFT process spans approximately 2 to 14 hours for models ranging from 0.5B to 7B in size. Pre-KD logits caching for a 7B model requires about 5 hours, whereas RL for the 3B model extends to approximately 3 days. The KD process for the 0.5B model is completed within about 2 hours, and training each pruned model requires roughly 3.5 days. Moreover, for the LLM prototype and LLM-as-a-Judge, generating data for every 10,000 entries takes approximately 1 hour.
Appendix F Detailed Analysis for Experimental Results for KD
As demonstrated in Table 6, it is clear that balancing FKL and RKL is crucial for our task. The function selection process by the model resembles a classification task, whereas generating parameter names and values for function calls is more akin to a generation task. The parameter $\mu$ in AKL is precisely used to adjust the effects between classification and generation. Experimental results demonstrate that at $\mu\leq 0.7$ and $\mu=1.0$ (equivalent to FKL), the performance with AKL is similar to or worse than RFT. Based on our findings where the average probability of the top-1 token from the teacher model exceeds 99%, we conclude that setting a lower $\mu$ value causes the effect of AKL to excessively shift more towards RKL. As described in the AKL paper, RKL prioritizes learning the tail of the probability distribution, which can impede rapid convergence in scenarios without requiring diversity. However, RKL can bring considerable benefits for the function classification part. Overall, AKL demonstrates greater potential compared to RFT, achieving performance improvements with an appropriately chosen $\mu$ value.
Appendix G Inference Details
The inference framework utilizes TensorRT-LLM[67] v0.16.0 with enabled optimizations including paged attention, prefix cache, CUDA graph size set to 1024, and max token number set to 32768 for inference acceleration.
Appendix H Pruning Details
H.1 Offline One-shot Pruning
Importance as Metric for Structured Pruning
For Depth Pruning, we adopt Layer Importance (LI) as the metric, which is based on the similarity between consecutive layers’ activation. Formally, we define the importance of the $i$-th layer as:
$LI_{i}=1-\mathbb{E}_{X,t}\frac{X_{i,t}^{T}X_{i+1,t}}{\|X_{i,t}\|_{2}\|X_{i+1,t}\|_{2}}$ (9)
where $X_{i}$ denotes the activation matrix of the $i$-th layer and $X_{i,t}$ denotes the $t$-th column (embedding vector) of $X_{i}$. After all layer activations $\{X_{i}\}_{i=1}^{L}$ through a single forward pass on calibration data are collected, the $N$ least important layers are removed, where $N$ is determined by the target pruning ratio. For Width Pruning, we classify width pruning into three dimensions based on channel dependencies: Embedding dimension, Head dimension and Intermediate dimension (as illustrated in Figure 8). For compact LLMs (e.g., Qwen2.5-0.5B with head size of 14), we exclude head dimension pruning due to its demonstrated sensitivity[37]. Instead, we employ an activation-based importance metric, Channel Importance(CI), for channel selection as follows:
$CI^{i}_{\text{embedding}}=\sqrt{\sum_{b,s}|\text{LN}(X)^{i}|^{2}},\quad CI^{i}_{\text{inter}}=\sqrt{\sum_{b,s}|x(W_{ffn1}^{i})^{T}|^{2}}$ (10)
where $CI^{i}_{\text{embedding}}$ denotes the importance score of the $i$-th embedding dimension, computed from the LayerNorm outputs of attention and FFN blocks, while $CI^{i}_{\text{inter}}$ denotes the importance of the $i$-th intermediate dimension, calculated from the $i$-th row of FFN’s gate and up projection matrices $W_{ffn1}^{i}$. The importance scores are first aggregated across batch and sequence dimensions through summation $\sum_{b,s}$, then normalized using L2 normalization within each subgroup to obtain global channel importance.
Figure 8: Width pruning dependency in the Qwen2.5 model.
Optimal Width Pruning Configuration
To ensure hardware-efficient matrix multiplication during inference, we set the pruning granularity to 64 for embedding dimension and 256 for the intermediate dimension. This approach results in multiple width-pruned model configurations, as show in Table 7. We calculate channel importance scores $CI^{i}_{\text{embedding}}$ and $CI^{j}_{\text{inter}}$ using a calibration dataset consisting of 1024 randomly sampled sequences from WikiText-2. For each configuration, we prune entire channel groups with the lowest CI scores and evaluate the resulting one-shot pruned models through WikiText-2 perplexity. Table 7 presents the optimal configurations for width pruning at ratio of 20% and 30%, identified as Qwen2.5-0.4B-Width-2 and Qwen2.5-0.35B-Width-3, respectively.
Table 7: Perplexity(PPL) evaluation on WikiText-2 comparing model architectures with one-shot width pruning at 20% and 30% pruning ratios.
Model
Hidden
Size
FFN
Inter Size
Non-Emb
Params(B)
Emb
Param(B)
Params
WikiText2
PPL
Qwen2.5-0.5B 896 4864 0.358 0.136 0.345 14.28
Qwen2.5-0.4B-Width-1 896 3328 0.259 0.136 0.395 349.89
Qwen2.5-0.4B-Width-2 832 3840 0.271 0.126 0.397 251.95
Qwen2.5-0.4B-Width-3 768 4352 0.278 0.117 0.395 354.74
Qwen2.5-0.4B-Width-4 706 4864 0.282 0.107 0.389 493.55
Qwen2.5-0.35B-Width-1 896 2560 0.209 0.136 0.345 1210.42
Qwen2.5-0.35B-Width-2 768 3584 0.236 0.117 0.351 1141.69
Qwen2.5-0.35B-Width-3 768 3584 0.236 0.117 0.353 1009.57
Qwen2.5-0.35B-Width-4 706 4096 0.242 0.107 0.349 2026.00
Qwen2.5-0.35B-Width-5 640 4864 0.256 0.097 0.353 2395.52
H.2 Recovery Training in Pruning
Recovery Training Dataset and Hyperparameters
We use the Infinity-Instruct-7M[78] dataset and apply AKL distillation[29] described in Section 3.2 for general instruction recovery training. Qwen2.5-7B-Instruct[11] serves as the teacher model, and we follow Minitron’s methodology[64] to perform teacher correction on the Infinity-Instruct-7M dataset. The Adam optimizer is used with a weight decay of 1e-4 in conjunction with a cosine learning rate schedule, decaying from 2e-5 to 2e-6. We conduct training over 2 epochs, utilizing a batch size of 32, with each batch packed to accommodate sequences of up to 7168 tokens.
Performance across Knowledge Assessment & Reasoning Tasks
Table 8 summarizes the performance of pruned models across various downstream tasks, including MMLU[79] and CMMLU[80] for knowledge assessment, alongside ARC-Challenge[81] and HellaSwag[82] for commonsense reasoning evaluation, and GSM8k[83] for logical reasoning. Our experiments reveal distinct recovery patterns between width and depth pruning across general tasks. At a 20% pruning rate, width pruning achieves 90.5% of the performance relative to the unpruned baseline, while depth pruning retains only 78.4%. When increasing the pruning intensity to 30%, width pruning preserves 83.8% of the original performance. In contrast, depth pruning suffers catastrophic performance degradation, particularly on the GSM8K task, where accuracy drops from 45.11 to a merely 5.61.
Table 8: Performance comparison of Qwen2.5 models after pruning. We report 5-shot performance on MMLU and CMMLU, 4-shot for GSM8k, zero-shot for ARC-Challenge, HellaSwag. * denotes that the baseline Qwen-2.5-0.5B utilizes the same instruct fine-tuning dataset for correction fine-tuning.
Model
MMLU
5-shot
CMMLU
5-shot
ARC-C
0-shot
HellaSwag
0-shot
GSM8k
4-shot
Avg.
Qwen-2.5-0.5B* 44.65 42.53 50.17 39.13 45.11 44.32
Qwen-2.5-0.4B-Width 38.63 35.89 42.27 35.55 48.60 40.12
Qwen-2.5-0.35B-Width 35.85 33.18 40.21 32.21 44.35 37.16
Qwen-2.5-0.4B-Depth 32.89 32.41 37.11 33.69 37.68 34.76
Qwen-2.5-0.35B-Depth 25.46 24.42 26.12 25.21 5.61 21.36
Ablation of Pruning Ratio on Domain-Specific Task
Our investigation reveals a nonlinear relationship between pruning intensity and model accuracy in domain-specific tasks. As shown in Table 9, both depth-wise and width-wise pruning at a 20% pruning rate result in a negligible <0.5% AR reduction. However, increasing the pruning rate to 30% leads to notable performance deterioration (>1.5% AR drop). These findings suggest that for small-scale models on domain-specific tasks, pruning rates exceeding 20% may significantly compromise accuracy.
Table 9: Results in different pruning strategies on domain-specific tasks.
Model AR
Qwen2.5-0.5B (KD) 94.85%
Qwen2.5-0.4B-Width (KD) 94.17%
Qwen2.5-0.35B-Width (KD) 93.71%
Qwen2.5-0.4B-Depth (KD) 94.20%
Qwen2.5-0.35B-Depth (KD) 93.75%
Appendix I FLOPs Calculation
For the calculation of FLOPs, we first define the notations symbolically:
  • $b$: the batch size
  • $h$: the hidden size
  • $s_{\text{t}}$: the total average number of prompt tokens
  • $s_{\text{u}}$: the average number of uncached tokens in prefill phase
  • $n_{\text{a}}$: the number of attention heads
  • $n_{\text{kv}}$: the number of separate key/value heads in GQA
  • $d_{\text{h}}$: the dimension of each head
  • $d_{\text{i}}$: the intermediate dimension of FFN layer
  • $v$: the vocabulary size
  • $l$: the number of model layers
According to [84], we calculate the floating-point operations (FLOPs) of the Qwen2.5 models, only considering matrix multiplications (GEMMs) without norms and activations:
  • Attention block operations:
    • Query/Key/Value transformations (GQA): $2bs_{\text{u}}hd_{\text{h}}(n_{\text{a}}+2n_{\text{kv}})$
    • Attention matrix computation and attention over values: $2bs_{\text{t}}s_{\text{u}}n_{\text{a}}d_{\text{h}}$
    • Post-attention linear projection: $2bs_{\text{u}}hn_{\text{a}}d_{\text{h}}$
  • Feed-forward network block operations:
    • Up/Gate/Down projections (each): $2bs_{\text{u}}hd_{\text{i}}$
  • LM head operations: $2bhv$
Thus, the total FLOPs for one inference with prefill cache is:
$\text{FLOPs}_{\ \text{total}}$ $=l\Bigg{[}4bs_{\text{u}}hd_{\text{h}}\left(n_{\text{a}}+n_{\text{kv}}\right)+2bs_{\text{u}}s_{\text{t}}d_{\text{h}}n_{\text{a}}+6bs_{\text{u}}hd_{\text{i}}\Bigg{]}+2bhv$
Workload and Computational Analysis
We conduct experiments on our datasets, in which prompts have an average length of 1792 tokens. By leveraging prefix caching to exploit shared prompt prefixes during inference, we reduce the average number of uncached tokens in the prefill phase to 128 and generate an average of 9 tokens per request. We calculate FLOPs under $s_{\text{u}}=128$ tokens in the prefill phase and under $s_{\text{u}}=1$ token for each step in the decoding phase, with detailed results presented in Table 1.
Appendix J Extensive Attempts on Generalization Enhancement
Table 10: In-domain and out-of-domain performance of Qwen2.5 models using different knowledge transfer strategies.
Model AR
AR
(Out-of-Domain)
BERT-based system 81.43% -
Qwen2.5-72B (ICL) 94.74% 94.63%
Qwen2.5-7B (RFT) 95.29% 90.86%
Qwen2.5-3B (RFT) 95.30% 91.02%
Qwen2.5-0.5B (RFT) 94.66% 90.62%
Qwen2.5-0.5B (KD) 94.85% 90.53%
Qwen2.5-7B (RL) 95.36% 94.80%
Qwen2.5-3B (RL) 94.68% 92.94%
Qwen2.5-0.5B (RL) 90.76% 85.21%
Knowledge Transfer using Qwen2.5-3B (RL)
Qwen2.5-0.5B (RFT*) 93.82% 89.13%
Qwen2.5-0.5B (RFT*+RL) 95.72% 90.09%
Qwen2.5-0.5B (KD*) 88.49% 84.14%
Qwen2.5-0.5B (KD*+RL) 95.55% 91.68%
Knowledge Transfer using Qwen2.5-7B (RL)
Qwen2.5-0.5B (RFT*) 95.29% 90.28%
Qwen2.5-0.5B (RFT*+RL) 95.94% 92.21%
Qwen2.5-0.5B (KD*) 95.32% 90.59%
Qwen2.5-0.5B (KD*+RL) 95.99% 92.34%
Figure 9: Response lengths of different models during RL training
Table 11: Response lengths of different models, represented as mean $\pm$ standard deviation.
Model Response Length
QwQ-32B (ICL) 316.8 $\pm$ 111.4
Qwen2.5-7B (RL) 86.4 $\pm$ 12.9
Qwen2.5-3B (RL) 61.7 $\pm$ 9.4
Qwen2.5-0.5B (KD*) 86.3 $\pm$ 12.6
Qwen2.5-0.5B (KD*+RL) 85.3 $\pm$ 10.8
Upon discovering that using RFT or KD alone leads to significant degradation in out-of-domain performance for super-tiny models, we conduct extensive experiments to enhance generalization capabilities, as illustrated in Appendix J. Initially, we prioritize Reinforcement Learning (RL) for this task. As expected, 7B RL-trained models achieve excellent results, surpassing the LLM Prototype and RFT-trained models. Subsequently, we attempt direct RL training on smaller models, such as 3B and 0.5B. The results are less favorable this time. While 3B RL-trained models exhibit considerable generalization capabilities, 0.5B RL-trained models show no improvement in generalization, underperforming even compared to 0.5B RFT-trained models. Through observations and analyses, we conclude that this is because models of such small size tend to generate low-quality Chain-of-Thought (CoT), as shown in Figure 9, leading to degraded output diversity, akin to scenarios where CoT is not used.
To address the challenge of 0.5B models’ inabilities to generate high-quality CoT, we propose a hybrid strategy. Given the favorable results from 7B RL-trained models, we decide to use 7B RL-trained models as our starting point. Then, we transfer out-of-domain capabilities and the abilities of generating high-quality CoT from 7B RL-trained models to 0.5B models via Rejection sampling Fine-Tuning (RFT) or knowledge distillation (KD), which are denoted as RFT* and KD*, respectively. Experimental results indicate that this knowledge transfer indeed outperforms direct RL training on 0.5B models in enhancing generalization and preserving the abilities of generating high-quality CoT.
So far, 0.5B KD* models posses the abilities to generate high-quality CoT, but in terms of out-of-domain Achievable Rate (AR), there is no considerable improvement compared to 0.5B RFT-trained models. Consequently, we conduct another round of RL training on 0.5B KD* models with data generated by the LLM prototype for enhancing generalization, similar to how we initially improve 7B models’ generalization with RL. Ultimately, experimental results present that after the second RL training, the 0.5B super-tiny models exhibit impressive improvement on model performance, with both an absolute increase of over 1% AR increase on in-domain and out-of-domain datasets compared to 0.5B RFT-trained models.
Table 12: Ablation of knowledge transfer strategies on Qwen2.5-0.5B models.
Model AR
AR
(Out-of-Domain)
Qwen2.5-0.5B (RFT*) 95.29% 90.28%
Qwen2.5-0.5B (KD*[FKL]) 95.08% 90.19%
Qwen2.5-0.5B (KD*[0.5FKL+0.5CE]) 95.32% 90.59%
Qwen2.5-0.5B (KD*[AKL]) 93.93% 88.37%
Qwen2.5-0.5B (KD*[0.5AKL+0.5CE]) 95.82% 90.85%
Qwen2.5-0.5B (RFT*+RL) 95.94% 92.21%
Qwen2.5-0.5B (KD*[FKL]+RL) 95.66% 92.21%
Qwen2.5-0.5B (KD*[0.5FKL+0.5CE]+RL) 95.99% 92.34%
Qwen2.5-0.5B (KD*[AKL]+RL) 95.85% 92.10%
Qwen2.5-0.5B (KD*[0.5AKL+0.5CE]+RL) 96.10% 92.03%
Ablation of Different Teacher Models
As illustrated in Appendix J, a teacher model with inherently superior capabilities is crucial for enhancing student model performance, particularly regarding out-of-domain capabilities. A better teacher model not only leads to better performance but also ensures greater stability in the student models’ results, driven by the generation of higher-quality CoT.
Response Lengths of Different Reasoning Models
As can be seen in Table 11, the average response lengths of our models are more than three times shorter than that of QwQ-32B. For direct RL training, longer response lengths appear to yield better results. However, this conclusion is not entirely valid. For instance, in the case of KD* models, there is almost no change in response lengths before and after RL, yet there is a significant difference in out-of-domain performance.
Ablation of Knowledge Transfer Strategies
During the knowledge transfer stage from 7B RL-trained Qwen2.5 models, we investigate the effectiveness of different knowledge transfer strategies, as illustrated in Appendix J. Apart from the KD*[AKL] strategy, which encounters a certain degree of underfitting, other experimental results demonstrate the robustness of different knowledge transfer strategies. Specifically, utilizing a loss with Cross Entropy (CE) is preferable to not incorporating CE during KD, as the high-quality CoT generated by 7B RL-trained models facilitates faster convergence with CE loss. Conversely, when the teacher model generates low-quality CoT, using CE alone, equivalent to RFT, can lead to overfitting, as evidenced in Appendix J, where the 3B RL-trained model as the teacher model exhibits poorer performance.
Appendix K Experiments on Public Datasets
Table 13: Performance of different models on BFCL and xlam-function-calling-60k datasets.
BFCL In-domain
Model Overall Non-live Live Acc
Qwen2.5-0.5B 55.14 62.29 47.98 55.88
Hammer2.0-0.5B [85] 59.31 67.00 51.62 /
Qwen2.5-72B 83.55 88.13 78.98 79.47
Qwen2.5-7B (RL) 81.85 86.27 77.42 78.55
Qwen2.5-0.5B (RL) 55.15 64.48 45.82 65.68
Qwen2.5-0.5B (RFT) 56.64 64.50 48.78 74.08
Qwen2.5-0.4B-Depth (RFT) 42.27 54.56 29.98 71.55
Qwen2.5-0.4B-Width (RFT) 42.60 56.17 29.02 72.98
Qwen2.5-0.5B-W8A16 (RFT) 56.42 64.42 48.41 74.38
Qwen2.5-0.5B-W4A16 (RFT) 49.54 58.52 40.56 72.23
Qwen2.5-0.5B-W8A8 (RFT) 55.68 62.88 48.48 74.06
Knowledge Transfer using Qwen2.5-7B (RL)
Qwen2.5-0.5B (KD*[FKL]) 59.44 69.06 49.81 72.97
Qwen2.5-0.5B (KD*[0.5FKL+0.5CE]) 54.28 62.00 46.56 66.55
Qwen2.5-0.5B (KD*[AKL]) 58.45 67.15 49.74 71.90
Qwen2.5-0.5B (KD*[0.5AKL+0.5CE]) 58.97 68.12 49.81 71.70
Qwen2.5-0.5B (KD*[FKL]+RL) 60.42 71.02 49.81 75.87
Qwen2.5-0.5B (KD*[0.5FKL+0.5CE]+RL) 59.84 70.00 49.67 77.02
Qwen2.5-0.5B (KD*[AKL]+RL) 60.50 69.56 51.44 76.38
Qwen2.5-0.5B (KD*[0.5AKL+0.5CE]+RL) 60.79 69.46 52.11 76.68
To validate the effectiveness of our proposed method, we conduct experiments on publicly available datasets. We randomly partition 90% of the xlam-function-calling-60k dataset [69] for training, with the remaining 10% serving as an in-domain test set. For the training set, instead of utilizing the provided labels, we employ Qwen2.5-72B-Instruct [11] to generate reference labels. Furthermore, we design a judge based on QwQ-32B [44] to assess the accuracy of the labels generated by the LLM. This process can be viewed as prototyping. Given that function calling is a standardized task, we use the official function call prompt template instead of developing a specialized agent for answer generation. The design of our judge is inspired by the Format Checker and Semantic Checker proposed in [69]. Through this procedure, we ultimately curate 50,201 valid query-tools-answers pairs from the initial 54k training samples for subsequent training.
Due to the complex function call formats within this dataset, we redesign the reward mechanism for Reinforcement Learning (RL). The reward is defined as:
$Reward=\alpha\cdot Reward_{format}+\beta\cdot Reward_{answer}$ (11)
where $Reward_{format}$ is 1 if and only if all the following conditions are met, and 0 otherwise: (1) The model’s output contains exactly one "<think>…</think>" block; (2) The function call within the answer can be parsed by the official Qwen prompt template and can be loaded as a JSON object; (3) The function name in the answer exists within the tools provided in the prompt, and the parameter types conform to their definitions, which is implemented using Pydantic[86].
For $Reward_{answer}$, we first check whether the number of functions in the model’s output is equivalent to that of the ground truth: If not, $Reward_{answer}$ is 0; Otherwise, $Reward_{answer}$ is defined as follows:
$Reward_{i}=\begin{cases}\frac{\lvert P_{i,pred}\cap P_{i,label}\rvert}{\lvert P_{i,pred}\cup P_{i,label}\rvert},&\text{if }\lvert P_{i,pred}\cup P_{i,label}\rvert>0\\ 1,&\text{otherwise}\end{cases}$ (12)
$Reward_{answer}=\frac{1}{N}\sum Reward_{i}$ (13)
where $N$ is the number of function calls, and $P_{i}$ denotes the set of parameters for the $i$-th function call. During training, we set $\alpha=1$ and $\beta=2$. The baseline variant of Reinforce++ [50] is selected as our reinforcement learning method.
We evaluate the trained model on both in-domain and out-of-domain dataset. For in-domain evaluation, we utilize the aforementioned test set with official answers as the ground truth. A test instance is considered correct only if the model’s output exactly matches the answer (i.e., identical function names and key-value pairs for arguments); all other cases are deemed incorrect. Accuracy is calculated based on this criterion. Additionally, we employ the Berkeley Function Call Leaderboard (BFCL)[70] for out-of-domain evaluation. The final experimental results are presented in Table 13.
The experimental results indicate that the 0.5B trained model achieves in-domain accuracy comparable to the Qwen2.5-72B model. Furthermore, it demonstrates a substantial improvement over Qwen2.5-0.5B-Instruct in out-of-domain scenarios. We also compare our method with Hammer2.0-0.5B[85], the previous SOTA 0.5B model, which is also fine-tuned from Qwen2.5-0.5B-Instruct8. Our approach achieves superior performance, demonstrating the effectiveness of our proposed method. In terms of quantization, experimental results demonstrate that applying W8A8 and W8A16 strategies results in negligible loss compared to the original model, while using the W4A16 strategy leads to significant loss in out-of-domain scenarios. Regarding pruning, all pruned models exhibit substantial losses in out-of-domain scenarios, yet own relatively minor loss in in-domain scenarios.